OpenStudy (anonymous):

please help me!!! I need help with the squeeze theream [x^2 sin(1/x)] lim x->0

OpenStudy (jamesj):

First, what are you aiming for .. in other words, what do you think the answer is and why, intuitively?

OpenStudy (anonymous):

its one i think but thats because someone told me

OpenStudy (anonymous):

its dne

OpenStudy (jamesj):

No. Have a look at the graph here:

OpenStudy (anonymous):

yeah ive seen it

OpenStudy (anonymous):

this is my teachers class that i'm in http://www.youtube.com/watch?v=9I2aamP-wB8&list=UU2el6NvSBUve08hVAqp4s6g&index=19&feature=plcp

OpenStudy (jamesj):

Now, from the graph, what intuitively is the answer? It's clearly not 1.

OpenStudy (anonymous):

it doesnt exist

OpenStudy (anonymous):

it nvr reaches 0

OpenStudy (jamesj):

No. It exists. You're not looking at the graph. It looks like it's zero, 0. And in fact that's the answer.

OpenStudy (jamesj):

Here's a hint for all. For all \( \theta \), \[ | \sin \theta | \leq 1 \]

OpenStudy (anonymous):

ok ell thanks for your help but its not making any sense and i'm just going to give up on it. but thanks

OpenStudy (anonymous):

I'm not quiting i'm just going to ask my teacher to explain it on paper. this word explaination is not doing it for me.